OpenStudy (anonymous):

Indicate in standard form the equation of the line passing through the given points. B(0, 0), C(4, -4)

OpenStudy (anonymous):

Standard form: y=mx+b, where m=slope and b=y-intercept. Using the equation m = (y2-y1)/(x2-x1) you can plug in the two points. Solving that you should get m = -1. After that, you can plug in one of the points you are given into the equation we now have y=(-1)x+b and solve for b. Using (4,-4) we get -4=(-1)(4)+b. b=0. Looking at the points you are given you can tell that the y intercept will be 0, as the orgin is one of the points given (0,0). So you end up getting y=-x as the equation of your line.